answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

What are the chances of not picking ace out of 52 cards?

There are four Aces in a standard deck of 52 cards. The probability, then, of not drawing an Ace is (52 - 4) in 52, or 48 in 52, or 12 in 13, or about 0.9231.

What are all the ways to combine 1 2 3 4 5 6 7 8 9 0 in 4 digits?

Digit 1 = 10 choices (0,1,2,3,4,5,6,7,8,9)

Digit 2 = 10 choices (0,1,2,3,4,5,6,7,8,9)

Digit 3 = 10 choices (0,1,2,3,4,5,6,7,8,9)

Digit 4 = 10 choices (0,1,2,3,4,5,6,7,8,9)

Number of combinations = 10 x 10 x 10 x 10 = 104 =10000

What does Brenna mean?

Brenna-Female version of Brennan. "Long Dark Hair" "Raven" "Raven" Comes from the Irish language.

1.Beautiful

2.Crazy

3.Funny

4.Hilarious

5.Intelligent

6.Wild

7.Young

A pair of fair dice is tossed and the output is the maximum number that comes on top what is the probability that the output is even?

when two dice are tossed , 6*6 = 36 outcomes are possible

a max output of 4 as defined can occur through

1-4 , 2-4 , 3-4 , 4-4 , 4-3 , 4-2 , & 4-1 i.e. 7 ways

so its probability is 7/36

the problem thus reduces to a binomial distribution

n = number of "trials" = 30

p = probability of "success" on a single trial = 7/36

q = 1 - p = probability of "failure" on a single trial = 29/36

the formula for k successes in n trials is given by

P[k] = nCk*p^k*q^(n-k)

thus reqd. probability = P[16] + P[17]

= 30C16*[7/36]^16*[29/36]^14 + 30C17*[7/36]^17*[29/36]^13

= 3.527/10^5

In blackjack what is the probability of being dealt three sevens?

The odds of being dealt three suited sevens in a six-deck game are approximately 63,000 to 1. [Source: Arnold Snyder, "The Big Book of Blackjack" p205]

The odds of being dealt three sevens (not suited) is somewhat worse than 500 to 1 (considering that casinos frequently offer the payout of 500 to 1 for players receiving three unsuited 7s in the "Super Sevens" side bet).

Is it possible not to die?

According to some religions, it's theoretically possible under extraordinary circumstances.

For example, Elijah and Enoch are stated not to have died, but to have been "taken" by God.

Also, it's implied in the New Testament in several places that all Christians who are still alive at the time of the Second Coming of Christ will not have to die.

Aside from that, though: No. Advances in medical technology could theoretically make it possible to extend a person's lifespan, but a non-zero cosmological constant implies that even that has a limit (though it's extremely far in the future) and that you'll die in the Big Rip if not before.

When do you add a probability?

Use of the word "or" is addition; e.g. probability of drawing out of a standard deck of cards a 9 or a 10 which is 4/52 + 4/52 = 8/52 or 2/13.

Drawing 3 cards that are all 9 from a deck of cards?

The probability of drawing 3 cards, all with the value of 9, from a standard 52 card deck, is ~0.018%.

Why the objects sink?

Objects sink when the density of the object is more than that of the water.

If you flipped three pennies what would be the odds that they all came out the same?

1 out of 4, or 25% Chance

HHH

HHT
HTH
HTT
THH
THT
TTH
TTT

There is only 3 coins, with 2 possible sides that they could land on, 8 possible outcomes, but with HHH and TTT both considered as the same side, that means there is 2 different out comes that are on the same side. There for, 2 out of 8, meaning 1 out of 4.

What is the probability to choose a prime number from 1 to 20?

In this problem, the total number of possibilities is 20, so n = 20.

The set of prime numbers from 1 to 20 = {2, 3, 5, 7, 11, 13, 17, 19}, so f = 8

Probability = f/n = 8/20 = 0.4.

You have a 2 in 5 chance of choosing a prime number from 1 to 20.

How does rolling dice apply to the real world?

Answer: The Heisenberg uncertainty principle is a fundamental part of quantum mechanics so chance is prerequisite to everything.

Which flag is half green and half red and also has a red crescent and a red star?

The flag of Algeria seems to be fairly close to your description. The related link shows the Algerian flag.

What cannot be a probability 0 -49 0.001 14 percent?

Probabilities can only range from 0 to 1. All of the options range from 0 to 1 (14% is the same as 0.14) except -49, so -49 is not a probability.

What does a probability of 1 mean?

This means the chances of this happening are close to, if not, 100%.

What is the probability of getting bit by a shark?

A shark will bite of its ben starved, if its provoked, or if it can sense that you are a good source of food for the moment.